# XDiag: Exact Diagonalization for Quantum Many-Body Systems

### v2

@@ 4.3 Tower of states
We perform a TOS analysis of the Heisenberg model on a square lattice with periodic boundary conditions. This model consists of spin $S=1/2$ sites at the vertices of a square lattice with nearest-neighbor antiferromagnetic Heisenberg couplings. The TOS analysis provides strong evidence for SSB in the thermodynamic limit, as the ground state of a finite system is completely symmetric. By plotting the spectrum obtained using XDiag vs. $S_{\text{tot}}(S_{\text{tot}}+1)$, we show that the low-energy spectrum of this model scales linearly with $S_{\text{tot}}(S_{\text{tot}}+1)$, a telltale sign of antiferromagnetic order.
@@
[…]
@@ 4.5 Time evolution
$|\psi_{0}\rangle=|\uparrow\ldots\uparrow\downarrow\ldots\downarrow\rangle,$ (35)
@@
where $h_{x}$ introduces quantum fluctuations and $h_{z}(t)=h_{z}^{\text{in}}+t/\tau_{Q}$ is the time-dependent longitudinal field. This system exhibits a first-order quantum phase transition (FOQPT) between two symmetry-broken ferromagnetic phases with opposite longitudinal magnetization. When the longitudinal field is slowly ramped from negative to positive values across the FOQPT, the system undergoes quantized bubble nucleation events [60]: sharp nonadiabatic transitions in the dynamics where local domains (or “bubbles”) of the new vacuum appear within the metastable phase.
@@ 6 Benchmarks
We benchmark the computation time of the key operation behind all main features implemented in XDiag: the execution of a single Lanczos iteration, consisting of a matrix-vector multiplication and a few vector operations. We investigate the performance of all possible Hilbert spaces supported by XDiag, namely the Spinhalf, Electron, and tJ blocks and, thus, consider three distinct models to perform the benchmarks:
@@
    $H=-t\sum_{\langle i,j\rangle\;\sigma}\left(c^{\dagger}_{i,\sigma}c_{j,\sigma}+\text{h.c.}\right)+U\sum_{i=0}^{N}n_{i,\uparrow}n_{i,\downarrow},$ (38)
    where $t$ is the hopping strength between first-neighboring sites, $c^{\dagger}_{i,\sigma},c_{i,\sigma}$ the creation and annihilation operator which creates (resp. destroys) a fermion on site $i$ with spin $\sigma=\uparrow,\downarrow$ and $U$ the strength of the on-site density-density interaction.
@@
    $H=-t\sum_{\langle i,j\rangle\;\sigma}P\left(c^{\dagger}_{i,\sigma}c_{j,\sigma}+\text{h.c.}\right)P+J\sum_{\langle i,j\rangle}\left(\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}-\dfrac{n_{i}n_{i+1}}{4}\right),$ (39)
@@
XDiag supports both multithreading using OpenMP and multiprocessing using MPI. As such, we provide benchmarks for the three models using both. We emphasize that in the current version of XDiag, lattice symmetries are not supported when using MPI parallelization. In all the following sections, the computation time always corresponds to the wall time and not to the CPU time.
@@ 6.1 Shared memory parallelization
[…]
Figure 3: Scaling of the computation time for a single Lanczos iteration as a function of the total number of threads used. The calculations were done for the Heisenberg $a)$, Hubbard $b)$ and t-J $c)$ models on a linear chain using different total numbers of spins/sites. The solid black line in each panel represents the slope corresponding to ideal linear scaling. In the calculation, we used both the U(1) and lattice symmetries.
@@
Block $N$ Symmetries Block Size Time (secs.)
@@ 6.2 Distributed memory parallelization
Figure 4: Scaling of the computation time for a single Lanczos iteration as a function of the total number of MPI processes used. The calculations were done for the Heisenberg $a)$, Hubbard $b)$, and t-J $c)$ models using different total numbers of spins/sites. The solid black line in each panel represents the slope corresponding to ideal linear scaling. In $a)$, the calculation was performed in the zero magnetization sector, in $b)$ it was done at half-filling for $N=20$, at $3/11$ for $N=22$ and $1/3$ for $N=24$.
[…]
Block $N$ Symmetries Block Size MPI processes Time (secs.)
@@ A.1 List of defined operator types
[…]
@@
Exchange A spin exchange interaction of the form $\frac{1}{2}(JS^{+}_{i}S^{-}_{j}+J^{*}S^{-}_{i}S^{+}_{j})$ 2 Spinhalf, tJ, Electron, SpinhalfDistributed, tJDistributed, ElectronDistributed
@@
tJSzSz An Ising interaction as encountered in the $t-J$ model of the form $S^{z}_{i}S^{z}_{j}-\frac{n_{i}n_{j}}{4}$ 2 tJ, tJDistributed
tJSdotS An Heisenberg interaction as encountered in the $t-J$ model of the form $\mathbf{S}_{i}\cdot\mathbf{S}_{j}-\frac{n_{i}n_{j}}{4}$ 2 tJ, tJDistributed
@@
A.2 Matrix type
C++
auto sx = arma::mat({{0, 1},{1, 0}});
auto op = Op("Matrix", 0, sx);
C++
auto sx = arma::mat({{0, 1},{1, 0}});
auto sz = arma::mat({{0.5, 0},{0, -0.5}});
arma::mat sxsz = arma::kron(sx, sz);
arma::mat sxszsxsz = arma::kron(sxsz, sxsz);
auto op_sxsz = Op("Matrix", {0, 1}, sxsz);
auto op_sxszsxsz = Op("Matrix", {0, 1, 2, 3}, sxsz);
Here, we have used the Kronecker product function kron.
A.3 Complex couplings
  1. 1.
  2. 2.
    $(tc^{\dagger}_{i\sigma}c_{j\sigma}+\textrm{h.c.})=(tc^{\dagger}_{i\sigma}c_{j\sigma}+t^{\star}c^{\dagger}_{j\sigma}c_{i\sigma})$
    $\frac{1}{2}(JS^{+}_{i}S^{-}_{j}+J^{\star}S^{-}_{i}S^{+}_{j})$

### v3

@@ 4.3 Tower-of-states
We perform a TOS analysis of the Heisenberg model on a square lattice with periodic boundary conditions. This model consists of spin $S=1/2$ sites at the vertices of a square lattice with nearest-neighbor antiferromagnetic Heisenberg couplings. The TOS analysis provides strong evidence for SSB in the thermodynamic limit, as the ground state of a finite system is completely symmetric. By plotting the spectrum obtained using XDiag vs. $S_{\text{tot}}(S_{\text{tot}}+1)$, defined as the eigenvalues of the total spin operator,
$\hat{\mathbf{S}}_{\mathrm{tot}}=\sum_{i}\hat{\mathbf{S}}_{i},\qquad\hat{\mathbf{S}}_{\mathrm{tot}}^{2}=\hat{\mathbf{S}}_{\mathrm{tot}}\cdot\hat{\mathbf{S}}_{\mathrm{tot}},$ (32)
we show that the low-energy spectrum of this model scales linearly with $S_{\text{tot}}(S_{\text{tot}}+1)$, a telltale sign of antiferromagnetic order.
@@
where $\langle\dots\rangle$ and $\langle\langle\dots\rangle\rangle$ denote sums over nearest and next-nearest neighbor sites, and $\langle\langle\langle\dots\rangle\rangle\rangle_{h}$ denotes a sum over third-nearest-neighbor sites [62, 36]. This model hosts a plethora of interesting phases for different parameter regimes. We focus on the parameter values $J_{2}/J_{1}=-1.0$ and $J_{3}/J_{1}=-2.0$, which were predicted [36] to realize a spin-nematic order with quadrupolar character. By plotting the low-energy spectrum, we show that odd-$S_{\text{tot}}$ sectors are not present in the low-energy TOS, indicating a quadrupolar spin-nematic phase.
@@ 4.5 Time evolution
$|{}_{0}\rangle=|\uparrow\ldots\uparrow\downarrow\ldots\downarrow\rangle,$ (38)
@@
where $h_{x}$ introduces quantum fluctuations and $h_{z}(t)=h_{z}^{\text{in}}+t/{}_{Q}$ is the time-dependent longitudinal field. This system exhibits a first-order quantum phase transition (FOQPT) between two symmetry-broken ferromagnetic phases with opposite longitudinal magnetization. When the longitudinal field is slowly ramped from negative to positive values across the FOQPT, the system undergoes quantized bubble nucleation events [64]: sharp nonadiabatic transitions in the dynamics where local domains (or “bubbles”) of the new vacuum appear within the metastable phase.
@@ 6 Benchmarks
We benchmark the computation time of the key operation behind all main features implemented in XDiag: the execution of a single Lanczos iteration, consisting of a matrix-vector multiplication and a few vector operations, and the generation of the Hamiltonian matrix as a sparse matrix in CSR format. We investigate the performance of all possible Hilbert spaces supported by XDiag, namely the Spinhalf, Electron, and tJ blocks and, thus, consider three distinct models to perform the benchmarks:
@@
[…]
@@
[…]
@@
XDiag supports both multithreading using OpenMP and multiprocessing using MPI. As such, we provide benchmarks for the three models using both. We emphasize that in the current version of XDiag, lattice symmetries are not supported when using MPI parallelization. Beyond parallel scalability, we also assess the efficiency against established libraries: we compare the serial performance of XDiag’s Lanczos routine with that of the ALPS library [13], and benchmark the sparse matrix generation against the QuSpin library [15, 16].
In all the following sections, the computation time always corresponds to the wall time and not to the CPU time.
Figure 3: Scaling of the computation time for a single Lanczos iteration as a function of the total number of threads used. The calculations were done for the Heisenberg $a)$, Hubbard $b)$ and t-J $c)$ models on a linear chain using different total numbers of spins/sites. The dashed black line in each panel represents the slope corresponding to ideal linear scaling. In the calculation, we used both the U(1) and lattice symmetries.
@@ 6.1 Shared memory parallelization
We show how the computation time of a single Lanczos step depends on the total number of OpenMP threads used for the shared memory parallelization as available. For the Spinhalf block we consider system sizes of $N=28$, $N=32$, $N=36$ and $N=40$ at zero magnetization, for the Electron block we consider system sizes of $N=16$, $N=18$, and $N=20$ at half-filling, and for the tJ block we consider system sizes of $N=28$, $N=30$, and $N=32$ in the presence of two holes. In addition to the U(1) symmetry, associated with the conservation of total magnetization in the Heisenberg model, or the conservation of total particle number in the Hubbard and t-J models, we also exploit the lattice translation symmetries to perform the calculations (see Section 3.10). Here, the Hamiltonian matrix elements are computed on-the-fly to maximize the memory available for storing the Lanczos vectors. The benchmarks presented were performed on a single node equipped with an Intel Xeon(R) Platinum $8360\text{Y}$ (Ice Lake) 2.4 GHz processor.
@@
Block NN Symmetries Block Size Time (secs.)
@@ 6.2 Distributed memory parallelization
Figure 4: Scaling of the computation time for a single Lanczos iteration as a function of the total number of MPI processes used. The calculations were done for the Heisenberg $a)$, Hubbard $b)$, and t-J $c)$ models using different total numbers of spins/sites. The dashed black line in each panel represents the slope corresponding to ideal linear scaling. In $a)$, the calculation was performed in the zero magnetization sector, in $b)$ it was done at half-filling for $N=20$, at $3/11$ for $N=22$ and $1/3$ for $N=24$.
XDiag also supports distributed-memory parallelization via MPI for the SpinhalfDistributed, ElectronDistributed, and tJDistributed blocks. We benchmark the computation time of a single Lanczos Iteration as a function of the number of MPI processes used. The calculations were performed on the Max-Planck-Gesellschaft Raven supercomputer using nodes also equipped with Intel Xeon(R) Platinum $8360\text{Y}$ (Ice Lake) 2.4 GHz processors, each featuring 72 cores. These nodes are interconnected with a 100 $\text{Gb}/\text{s}$ HDR100 network. In these benchmarks, lattice symmetries are not used, only U(1) number conservation symmetries. We consider the case of $N=36$ and $N=40$ for the Heisenberg model at zero magnetization. For the Hubbard model, we consider $N=20$ at half-filling, and $N=22$ and $N=24$ with $n_{\uparrow}=n_{\downarrow}=8$ electrons. The $t$-$J$ model is studied on $N=28$ and $N=32$ sites with $2$ holes at zero magnetization. We observe that for all systems considered, the computation time scales almost inversely proportional to the total number of MPI processes, as shown in Fig, 4. Moreover, we provide, in the table 2, the actual computational time required for each Lanczos iteration across the different models considered. We conclude that the MPI parallelization exhibits almost linear strong scaling up to several thousand processes.
Block NN Symmetries Block Size MPI processes Time (secs.)
@@
6.3 Sparse Matrix Generation
Figure 5: Computational time for the generation of the Hamiltonian matrix in the sparse CSR matrix format as a function of the total number of threads used. The calculations were done for the Heisenberg $a)$, Hubbard $b)$, and t-J $c)$ models using different total numbers of spins/sites. The dashed black line in each panel represents the slope corresponding to ideal linear scaling.
XDiag supports the creation of several common sparse matrix formats. In particular, the CSR format can be used directly by the library’s iterative solvers. In this section, we analyze the computation time required to generate the CSR matrix and evaluate its scalability with respect to the number of OpenMP threads. We consider the Hamiltonian for three distinct models with the following system sizes: (i) Spinhalf with $N=24$, $N=28$, and $N=32$; (ii) Electron with $N=10$, $N=12$, $N=14$, and $N=16$; and (iii) tJ with $N=20$, $N=22$, $N=24$, and $N=26$. In all calculations, we exploit again both lattice translation symmetries and the relevant U(1) symmetry. Specifically, we consider the $S^{z}=0$ sector for the Heisenberg model, half-filling block for the Hubbard model, and the two-hole sector with $S^{z}=0$ for the t-J model. These benchmarks were performed on a single node equipped with an Intel Xeon Platinum 8360Y (Ice Lake) 2.4 GHz processor.
Block NN Symmetries Block Size Time (secs.)
Spinhalf 32 U(1) and Lattice $9.4\cdot 10^{6}$ $(6.40\pm 4)\cdot 10^{3}$
Spinhalf 28 U(1) and Lattice $7.2\cdot 10^{5}$ $(4.6\pm 2)\cdot 10^{2}$
Spinhalf 24 U(1) and Lattice $5.6\cdot 10^{4}$ $(8.8\pm 2)\cdot 10$
Electron 16 U(1) and Lattice $1.0\cdot 10^{7}$ $(2.47\pm 0.02)\cdot 10^{3}$
Electron 14 U(1) and Lattice $8.4\cdot 10^{5}$ $(1.99\pm 0.05)\cdot 10^{2}$
Electron 12 U(1) and Lattice $7.1\cdot 10^{4}$ $(2.2\pm 0.1)\cdot 10$
Electron 10 U(1) and Lattice $6.4\cdot 10^{3}$ $(1.32\pm 0.07)\cdot 10$
tJ 26 U(1) and Lattice $3.4\cdot 10^{7}$ $(1.845\pm 0.008)\cdot 10^{4}$
tJ 24 U(1) and Lattice $8.1\cdot 10^{6}$ $(4.070\pm 0.005)\cdot 10^{3}$
tJ 22 U(1) and Lattice $1.9\cdot 10^{6}$ $(9.5\pm 0.1)\cdot 10^{2}$
tJ 20 U(1) and Lattice $4.6\cdot 10^{5}$ $(2.22\pm 0.07)\cdot 10$
Table 3: Computation times for the generation of the Hamiltonian matrix in the CSR format for each block type when $n_{\text{threads}}=64$.
Fig. 5 illustrates the scaling of computational time as a function of the number of OpenMP threads. For larger system sizes, we observe a significant decrease in the computational time as the number of threads increases. Notably, for the Heisenberg model $a)$, the computational time scales almost linearly with the inverse of the thread number up to $n_{\rm threads}=8$. In contrast, the Hubbard $b)$ and tJt-J $c)$ models exhibit a more pronounced deviation from ideal linear scaling. The actual computational time for $n_{\rm threads}=64$ is presented in Table 3 for the different system sizes and models.
6.4 Comparison with Other ED Solvers
Finally, we assess XDiag’s performance relative to existing ED libraries. We compare the serial performance of the Lanczos routine against the one in the ALPS library [13] and benchmark the sparse matrix generation against the QuSpin library [15, 16]. For both benchmarks, we restrict the comparison to serial performance, as these operations are not parallelized in ALPS or QuSpin.
Panels $a)$ and $b)$ of Fig. 6 illustrate the performance of QuSpin and XDiag in generating the Hamiltonian matrix in CSR format. Panel $a)$ considers the Hamiltonian matrix in the zero magnetization, zero momentum, and even inversion symmetry sectors, while panel $b)$ corresponds to the half-filling, zero momentum, and even inversion symmetry sectors. We observe that the XDiag implementation consistently outperforms QuSpin, with computation times lower by an order of magnitude. Furthermore, XDiag allows for multithreaded parallelization, which can reduce the computational time even further.
In panel $c)$, we compare XDiag’s Lanczos performance against that of the ALPS library. The calculations represent the computational time required for five Lanczos iterations on a Heisenberg chain in the zero magnetization and zero momentum sectors (point-group symmetries were not used as they are not currently supported by ALPS). We observe that XDiag performs better than ALPS across all system sizes considered.
Figure 6: Computational efficiency of XDiag compared to QuSpin and ALPS. Panels $a)$ and $b)$ illustrate the scaling of computational time for generating the Hamiltonian matrix (CSR format) for different system sizes, contrasting XDiag with the QuSpin library. Panel $c)$ benchmarks the Lanczos iteration performance of XDiag against the ALPS library, using a Heisenberg chain in the zero magnetization and momentum sectors across a range of system sizes.
@@ A.1 List of defined operator types
Hopup A hopping term for $\uparrow$ spins of the form ${\color[rgb]{1,0,0}-}(tc^{\dagger}_{i\uparrow}c_{j\uparrow}+\textrm{h.c.})$ 2 tJ, Electron, tJDistributed, ElectronDistributed
Hopdn A hopping term for $\downarrow$ spins of the form ${\color[rgb]{1,0,0}-}(tc^{\dagger}_{i\downarrow}c_{j\downarrow}+\textrm{h.c.})$ 2 tJ, Electron, tJDistributed, ElectronDistributed
@@
[…]
@@
tJSzSz An Ising interaction as encountered in the tJt-J model of the form $S^{z}_{i}S^{z}_{j}-\frac{n_{i}n_{j}}{4}$ 2 tJ, tJDistributed
tJSdotS An Heisenberg interaction as encountered in the tJt-J model of the form $\mathbf{S}_{i}\cdot\mathbf{S}_{j}-\frac{n_{i}n_{j}}{4}$ 2 tJ, tJDistributed